OpenStudy (anonymous):

S=So+Vo(t) S=So+Vo(t)+[a(t)^2]/2 V=Vo+a(t) V^2=Vo^2 + 2a(S-So)

OpenStudy (anonymous):

what is your question ?

OpenStudy (anonymous):

\[v_{average} = (v_f+v_0)/2\] \[v_f = v_0+at\] \[\Delta s = v_{average}t \rightarrow t*(v_f+v_0)/2 \rightarrow t*(v_0+at+v_0)/2\] \[\Delta s = v_{average}t \rightarrow t*(v_0+at+v_0)/2 \rightarrow v_0t+(1/2)at^2\] \[(v_f)^2 =( v_0+at)^2 \rightarrow v_0^2+2v_0at+(at)^2\rightarrow v_0^2+2a(v_0t+(1/2)at^2) \] \[v_f^2 = v_0^2+2a \Delta s\]